Hello,
I'm having a hard time figuring this out.
The side of an equilateral triangle is 4 inches shorter than the side of a square. The perimeter of the square is 38 inches more than the perimeter of the triangle. Find the length of the side of the square.
Thank you!

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

The side of an equilateral triangle is 4 inches shorter than the side of a square. The perimeter of the square is 38 inches more than the perimeter of the triangle. Find the length of the side of the square.
-----------
Let a side of the square be "s".
Its perimeter = "4s".
-----
A side of the equilateral triangle is "s-4"
Its perimeter = 3(s-4) = 3s-12
------------------------------------
Equation:
Perimeter of square - Perimeter of Equilater Triangle = 38
4s - (3s-12) = 38
s = 26 inches (length of a side of the square)
=================================================
Cheers,
Stan H.
===============

Answer by solver91311(24713)



The perimeter of a square is given by multiplying 4 times the measure of one of the sides. The perimeter of an equilateral triangle is given by multiplying 3 times the measure of one of the sides. Your triangle has sides that are 4 less than the measure of the side of the square, so if we let represent the measure of the side of the square, then represents the measure of the side of the triangle. Therefore:



Just solve for

John